OpenStudy (anonymous):

At a local coffee shop, an employee noticed that for every 24 customers, 8 ordered a pastry. If the coffee shop plans to have 1,440 customers this month, what is a reasonable prediction of the number of pastries that will be ordered?

OpenStudy (anonymous):

530 480 720 240

OpenStudy (anonymous):

8/24=x/1440 x=(1440x8)/24 x=480